OpenStudy (anonymous):

2x + 3y = 1,200 In slope intercept form? Medal will be given!

OpenStudy (anonymous):

y = (-2/3)x + 400 answer i believe

OpenStudy (anonymous):

Thank you! Do you have any work? :)

OpenStudy (anonymous):

@iPwnBunnies can you help? :)

OpenStudy (anonymous):

yes i do @MiaKitty1

OpenStudy (anonymous):

You do? :) Can you show me what u did? :) I appreciate the help you are giving me! :)

OpenStudy (anonymous):

@phi could you help me?

OpenStudy (ipwnbunnies):

You got this @tyree20 ?

OpenStudy (anonymous):

y – y1 = m(x – x1) I know this is the formula, I don't understand how to put it in. :(

OpenStudy (anonymous):

2x + 3y = 1200 -2x + 2x + 3y = -2x + 1200 0 + 3y = -2x + 1200 3y = -2x + 1200 (1/3)3y = (1/3)(-2x + 1200= ans:y = (-2/3)x + 400 @MiaKitty1

OpenStudy (anonymous):

That's it? Thanks for ur help! :)

OpenStudy (anonymous):

Do you know how to graph based on slope intercept form?

OpenStudy (tfht):

y=mx+b b=the y coordinate and m=slope. so you would just figure out what is in y's spot then put that at (0,y) then use the slope to go up so many and over so many

OpenStudy (anonymous):

Oh, okay, thanks sooo much! :D

OpenStudy (tfht):

so like if it was y=4/3x+7, you would start at (0,7) then go up 4 and to the right 3

OpenStudy (anonymous):

Okay, that makes sense, thanks for the help, I appreciate it! :D

OpenStudy (tfht):

yep!